What is the molality of a solution of 2.50 g H3PO4 dissolved in 240 g water?

1
Expert's answer
2022-03-10T10:44:42-0500

b(molality) = n(solute)/m(solvent);

M(H3PO4) = 98 g/mol;

n(H3PO4) = m/M = 2.5/98 = 0.0255 moles;

b(H3PO4) = n(H3PO4)/m(H2O) = 0.0255/0.24 = 0.11 mol/kg.

Answer: 0.11 mol/kg.